Through: (2,-5), parallel to y=4x-5
Answer by jim_thompson5910(35256)


y = mx+b

y = 4x+b

-5 = 4(2)+b

-5 = 8+b

-5-8 = b

-13 = b

b = -13

So the equation of the parallel line is y = 4x - 13